INAF LBC+LUCIFER queue observing, 2010 December 04 - 05

Observer: K.Boutsia, M. Fumana,V. Testa
Telescope Operator: S. Allanson
Support Astronomer: M. Pedani
Instrument: LBC + LUCIFER

Summary

Clouds at the beginning of the night. We try to start with LBC anyway.
The clouds persist and we loose 30min because of weather.
The first half of the night had good seeing but a lot of clouds.
During the second half of the night the sky was almost clear and we had a very good seeing <0.5''

Data Summary

Project Objects Filter ExpTime Notes
Clementini AndXXI - AndXIX B-bessel, V-bessel 10min + 10min too much extinction, out of specs
Pentericci cluster2HK 200JH OrderSep 4hr ARC+FLAT , passing clouds, seeing is oscillating from 0.6'' to 1.5''
Palazzi GRB070306 210zJHK + J 2hr ARC+FLAT+Tell1+Tell2, seeing<0.5''
Palazzi GRB070306 210zJHK + H 0.5hr ARC+FLAT+Tell2, seeing<0.5''

Details

00:07 - REBOOT LBC

00:25 - Connect LBC@LBT, cyrrhus everywhere
  • Wait for Sunset
  • 5Bias_Bino

00:40 -Sky Flats in BV
  • BV rot0 - 1.5sec
  • BV rot180 - 8sec

00:55 - Slew to pointing star WT10_351
  • ptautoadjust - mirradjust

01:06 - Slew to Clementini AndXXI - dofpia
  • NO GOOD PUPILS FOUND
  • give manually Z8=-2000 and Z7=-2000 to both sides and try again
  • dofpia,/First,/X2
  • still no good pupils - probably because of clouds
  • we keep repeating dofpia

01:50 - observing target Clementini - M31_AndXXI_420s_16
  • seeing ~1.0''
  • background too high - too much extinction - loosing 3.5mags
  • OB aborted after second exposure - we also lost guide star

02:07 - SWITCH TO LUCIFER

02:27 - correcting pointing near Pentericci cluster2 field

02:34 - Slew to Pentericci cluster2HK for MOS acquisition 0001
  • preset had to be aborted because guide star not well centered
  • restart AGW and GCS

02:42 - Slew to Pentericci cluster2HK for MOS acquisition 0002-0005

02:57 - Science spectra for cluster2HK 0006- 0012
  • We lost guide star - probably because of clouds
  • Preset canceled during the 8th offset

03:36 - repeat MOS acquisition Pentericci cluster2HK 0013-0014
  • it is taking too long to collimate - we switch off waiting for collimation
  • guide star still very faint - probably clouds passing by
  • stars too faint to do MOS acquisition

03:52 - we repeat MOS acquisition increasing exposure time 0015-0018
  • we could not see the pivot stars through the slit this time - probably because too faint
  • the offsets calculated by MOS ACQ are the same as last time and we decide to go on with the sci script

04:05 - Science spectra for cluster2HK 0019-0029
  • guide star very variable because of passing clouds
  • average seeing ~0.7''
  • lost preset because there is no guide star
  • OB interrupted after the 11th offset - 55minutes executed

05:05 - No stars visible outside! We wait

05:33 - Looks better - we repeat MOS acquisition 0030-0033

05:45 - Science spectra for cluster2HK for 2nd hr 0034-0045
  • when guiding star visible and in focus the seeing is ~0.6''

06:52 - Science spectra for cluster2HK for 3rd hr 0046-0057
  • giving an offset Y=-1.5 to clear overall offsets before launching science script
  • Seeing is oscillating a lot from 0.6 to 1.3

07:57 - Science spectra for cluster2HK for 4th hr 0058-0063
  • still very variable seeing
  • OB interrupted after 6th offset
  • this way we made up for the 6 missing offsets earlier and finished 4 hours

08:30 - Calibration Unit IN
  • ARC 0064-0067
  • FLAT 0068-0071

08:45 - Slew to GRB070306 field - acquisition Telluric1 for Palazzi GRB070306 0072-0077
  • It took some time to change the mask - more than usual

09:06 - Spectra for Telluric1 in J band
  • the first images were too bright, probably saturated 0078-0083
  • we have decreased the time from 30DITx2NDIT to 15DITx2NDIT 0084-0089

09:18 - Acquisition for GRB070306 (Palazzi) 0090-0099
  • we had trouble with the shift image command
  • it was giving too big offsets - we cannot see the object in the slit
  • we repeated preset and we have shifted manually

09:58 - Science spectra for GRB070306 in J 0100-0109
  • seeing <0.5''

12:02 - Calibration unit IN
  • ARC J 0110-0115
  • FLAT J 0116-0139

12:11 - Calib Unit OUT - Slew to GRB070396 Telluric2 for acquisition 0140-0143

12:19 - Spectra for Telluric2 in J band 0144-0149

12:23 - Acquisition for GRB070306 (Palazzi) 0150-0153
  • this time shift image worked properly

12:33 - Science spectra for GRB070306 in H 0154-0156
  • OB aborted after 1st offset because sky too bright, 30min executed

13:06 - Calibration Unit IN
  • ARC H 0157-0161
  • FLAT H 0162-0180

13:15 - Calibration Unit OUT - Slew to GRB070396 Telluric2 for acquisition 0181-0183

13:22 - Spectra for Telluric2 in H band
  • the first images were too bright, probably saturated 0184-0185
  • we have decreased the time from 30DITx2NDIT to 15DITx2NDIT 0186-0203

13:35 - END OF NIGHT

-- VincenzoTesta - 05 Dec 2010
Topic revision: r3 - 15 Mar 2012, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback